OpenStudy (anonymous):

An inequality is shown. 12 + 11/6x ≤ 5 + 3x What is the solution set

OpenStudy (jdoe0001):

http://openstudy.com/study#/updates/56c6633ae4b0d0b671577ea0 so.. if you were to solve \(\bf 12+\cfrac{11}{6x}= 5+3x\) for "x" what would it look like?

OpenStudy (anonymous):

x ≤ 6/13? The x next to the 6 should be next to the entire fraction.

OpenStudy (jdoe0001):

in short, you pretty much, simplify the iniequality, in the same way you would, an EQUALity, or EQUATion there's only a minor difference, is all, but is basically about the same

OpenStudy (jdoe0001):

x ≤ 6/13? <-- well, that's not an equation though did you mean x = 6/13?

OpenStudy (anonymous):

If you solve for x with it, the result is x ≥ 6.

OpenStudy (jdoe0001):

well... work it as an equation, that is, as \(\bf 12+\cfrac{11}{6x}= 5+3x\) see what you get for "x" :)

OpenStudy (jdoe0001):

ohhh hmm that is \(\bf 12+\cfrac{11}{6}x= 5+3x\)

OpenStudy (jdoe0001):

hint: multiply both sides by 6

OpenStudy (anonymous):

So if I'm correct, I am supposed to multiply the 12 by 6. I just don't remember how to multiply the other side.

jhonyy9 (jhonyy9):

so the other side will be 6(5+3x) yes ?

OpenStudy (anonymous):

So 6 x 8x?

jhonyy9 (jhonyy9):

no 6*5 +6*3x = ?

OpenStudy (anonymous):

18x + 30? Which would continue to 48x?

jhonyy9 (jhonyy9):

no because you not can assuming aple with melon - yes ?

jhonyy9 (jhonyy9):

18x will be the term with unknowed x and 30 an integer

jhonyy9 (jhonyy9):

so this will be on the right side but dont forget the left side what will be ?

OpenStudy (anonymous):

6 x 12?

jhonyy9 (jhonyy9):

yes 6*12 +11x ok.?

OpenStudy (anonymous):

it would lead up to 11 x+72?

jhonyy9 (jhonyy9):

so these will be on the left side but there are again what you have calculed above for the right side can you rewrite them now like an equalitie ?

OpenStudy (anonymous):

Would it be 11 x+72 > 18x + 30?

jhonyy9 (jhonyy9):

where you posted this exercise the leftside is less or equale by right side

jhonyy9 (jhonyy9):

so will be 11x+72 <= 18x+30

OpenStudy (anonymous):

Ah, Okay.

jhonyy9 (jhonyy9):

can you solve it now ?

jhonyy9 (jhonyy9):

subtract from both sides 11x so like a first step

jhonyy9 (jhonyy9):

what will get ?

jhonyy9 (jhonyy9):

11x -11x +72 <= 18x-11x +30 can you continue ?

OpenStudy (anonymous):

72<=7 x+30?

jhonyy9 (jhonyy9):

yes so now what will be the next step ?

jhonyy9 (jhonyy9):

subtract from both sides 30 so what will get ?

OpenStudy (anonymous):

30 - 72<= 30 - 7 x+30?

jhonyy9 (jhonyy9):

no subtract from both sides mean 72-30 <= 7x+30-30

OpenStudy (anonymous):

So 42 <= 7x? Since the +30 essentially gets removed by the -30?

jhonyy9 (jhonyy9):

yes - right so sure so now have got 42 <= 7x how you get the x ?

OpenStudy (anonymous):

Not entirely sure to be honest.

jhonyy9 (jhonyy9):

can you solve it now for x ?

OpenStudy (anonymous):

x>=6

jhonyy9 (jhonyy9):

yes so good - this is the right answer ok. ?

OpenStudy (anonymous):

Thanks.

jhonyy9 (jhonyy9):

was my pleasure any time good luck bye